OpenStudy (anonymous):

Which of the following statements is true? a. Any number raised to the zero power equals zero. b. Any number raised to a negative power equals a negative number. c. Negative numbers can be written in scientific notation. d. The product of two powers with the same base equals the base raised to the sum of the exponents.

OpenStudy (anonymous):

c or d

OpenStudy (anonymous):

a, i think

OpenStudy (anonymous):

anything to the 0 is 1

OpenStudy (anonymous):

no it = 1

OpenStudy (aroub):

no its not a

OpenStudy (anonymous):

it c or d

OpenStudy (anonymous):

it is not A.

OpenStudy (anonymous):

a is false x^0=1 for x=\=0, b is false x^(-y)=1/(x^y),

OpenStudy (anonymous):

yes its not a

OpenStudy (anonymous):

because it becomes 0.???

OpenStudy (anonymous):

any number raised to zero is 1, not zero (other than 0 itself, which is undefined)

OpenStudy (anonymous):

-100 in scientific notation -1.0 * 10^2

OpenStudy (anonymous):

a*a*a = a^3 a*a = a^2 a^1 = 1*a^1 = a a^0 = 1*a^0 = 1

OpenStudy (anonymous):

C is true, but im not sure of D, too tired

OpenStudy (anonymous):

a is not true as 0^0 is undefined